# SPIN-Bench: How Well Do LLMs Plan Strategically and Reason Socially?

### v1

@@ A.1 Classification and description of PDDL Domains
A.2 Two-player Competitive Games
@@
A.3 Multi-player Cooperative Game - Hanabi
@@
In our Hanabi experiment, we tested 9 LLMs in total: o1, o1-mini, o3-mini, GPT-4o, GPT-4-turbo, Claude 3.5 Sonnet, Claude 3.5 Haiku, DeepSeek R1, Llama-3.3-70b. However, Llama-3.3-70b failed to play this game, so we only reported the above 8 models’ results. For each model, we have four settings: allowing the player number to range from 2 to 5. For example, model o1 with a player number of 3 means there are 3 separate o1 models playing the game. And we use the final score when the game ends as the score of each game setting.
A.4 Strategic Game - Diplomacy
A.4.1 Game Introduction
@@ A.4.2 Evaluation Metrics
For each negotiation phase in a complete game trajectory (20 years), we provide the annotation prompt to o1 model, gathering all the results from its analysis, and compute the statistics. An example of one negotiation phase and o1 model’s annotation is shown in Appendix C.2.
Appendix B Supplementary Experiments
B.1 Two-player Competitive Games
Model Leaderboard
| Model | Tic Tac Toe | Connect Four | Chess |
| Human | 1415 | 1692 | - |
| o1-preview | 1263 | 1377 | 1395 |
| o1-mini | 1205 | 1083 | 1247 |
| Claude 3.5 sonnet | 1138 | 942 | 1196 |
| GPT-4o | 1025 | 980 | 1255 |
| Claude 3.5 haiku | 983 | 907 | 1041 |
| Qwen2.5:72b | 944 | 877 | 1150 |
| Llama3.3 | 944 | 878 | 1146 |
| Llama3.1:70b | 936 | 829 | 1164 |
| GPT-4o-mini | 912 | 925 | 842 |
| GPT-4-turbo | 857 | 1007 | 1265 |
| Llama3:70b-instruct | 856 | 993 | 1155 |
| Mistral:7b | 855 | 800 | 243 |
| Llama3.2:3b | 836 | 821 | 421 |
| GPT-3.5-turbo | 829 | 890 | 480 |
Table 6 presents the internal Elo ratings of various LLMs across three competitive games: Tic Tac Toe, Connect Four, and Chess. It is important to note that these ratings are internal metrics and do not directly reflect performance among human players. Among evaluated models, o1-preview and o1-mini lead in this category, demonstrating strongest strategic planning capabilities over the remaining models, further indicating the advantages of test time scaling in strategic games. Additionally, Claude 3.5 Sonnet, GPT-4o, and GPT-4-turbo maintain relatively high Elo ratings, indicating a superior ability to handle complex strategies compared to their counterparts. In contrast, models such as Mixtral, Llama3.2 perform significantly worse in Chess reflecting the increased demands that games with larger action and state spaces impose on a model’s planning and reasoning capabilities. However, for Tic Tac Toe and Connect Four, human outperforms all LLMs by a substantial margin. This result underscores the current limitations of LLMs in strategic reasoning, highlighting the significant gap between AI and human-level planning capabilities.
@@ B.1 Two-player Competitive Games
| o1-preview | 10.0 | 100.0 | 100.0 |
| o1 | 30.0 | 100.0 | 100.0 |
| GPT-4-turbo | 40.0 | 100.0 | 100.0 |
| Claude 3.5 Sonnet | 40.0 | 100.0 | 100.0 |
| o1-mini | 50.0 | 100.0 | 100.0 |
| Claude 3.5 Haiku | 50.0 | 100.0 | 100.0 |
| Llama3:70b | 80.0 | 100.0 | 100.0 |
| GPT-4o-mini | 80.0 | 100.0 | 100.0 |
| Llama3.1:70b | 80.0 | 100.0 | 100.0 |
| o3-mini | 80.0 | 100.0 | 100.0 |
| Qwen2.5:72b | 90.0 | 100.0 | 100.0 |
| Llama3.2:3b | 90.0 | 100.0 | 100.0 |
| GPT-3.5-turbo | 90.0 | 100.0 | 100.0 |
| DeepSeek-R1 | 90.0 | 100.0 | 100.0 |
| Llama3.3:70b | 100.0 | 100.0 | 100.0 |
| GPT-4o | 100.0 | 100.0 | 100.0 |
B.2 Diplomacy
B.2.1 Basic Skill Evaluation
@@
B.2.2 Multi-agent Experiment
@@
Appendix C Prompt Design
C.1 Cooperative Game: Hanabi
@@
C.2 Strategic Game: Diplomacy
@@
Example Annotation by o1:

### v3

@@ B.1 Classification and description of PDDL Domains
B.2 Two-player Competitive Games
@@
| Game | Game Tree Complexity | Branching Factor | State Space Complexity |
| (as log to base 10) | (as log to base 10) | (as log to base 10) | |
| Tic Tac Toe | 5 | 4 | 3 |
| Connect Four | 21 | 4 | 13 |
| Chess | 123 | 35 | 44 |
@@
B.3 Multi-player Cooperative Game - Hanabi
@@
(Note: The following game introduction is excerpted from the Hanabi Wikipedia Page)
@@
In our Hanabi experiment, we tested 9 LLMs in total: o1, o1-mini, o3-mini, GPT-4o, GPT-4-turbo, Claude 3.5 Sonnet, Claude 3.5 Haiku, DeepSeek R1, Llama-3.3-70b. For each model, we have four settings: allowing the player number to range from 2 to 5. For example, model o1 with a player number of 3 means there are 3 separate o1 models playing the game. And we use the final score when the game ends as the score of each game setting.
B.4 Strategic Game - Diplomacy
B.4.1 Game Introduction
@@ B.4.2 Evaluation Metrics
For each negotiation phase in a complete game trajectory (20 years), we provide the annotation prompt to o1 model, gathering all the results from its analysis, and compute the statistics. An example of one negotiation phase and o1 model’s annotation is shown in Appendix D.3.
Appendix C Supplementary Experiments
C.1 Two-player Competitive Games
@@
| o1-preview | 90.0 | 0.0 | 0.0 |
| o1 | 70.0 | 0.0 | 0.0 |
| GPT-4-turbo | 60.0 | 0.0 | 0.0 |
| Claude 3.5 Sonnet | 60.0 | 0.0 | 0.0 |
| o1-mini | 50.0 | 0.0 | 0.0 |
| Claude 3.5 Haiku | 50.0 | 0.0 | 0.0 |
| Llama3:70b | 20.0 | 0.0 | 0.0 |
| GPT-4o-mini | 20.0 | 0.0 | 0.0 |
| Llama3.1:70b | 20.0 | 0.0 | 0.0 |
| o3-mini | 20.0 | 0.0 | 0.0 |
| Qwen2.5:72b | 10.0 | 0.0 | 0.0 |
| Llama3.2:3b | 10.0 | 0.0 | 0.0 |
| GPT-3.5-turbo | 10.0 | 0.0 | 0.0 |
| DeepSeek-R1 | 10.0 | 0.0 | 0.0 |
| Llama3.3:70b | 0.0 | 0.0 | 0.0 |
| GPT-4o | 0.0 | 0.0 | 0.0 |
C.2 Diplomacy
C.2.1 Basic Skill Evaluation
@@
C.2.2 Multi-agent Experiment
Table 10 shows the full results for multi-agent Diplomacy experiments.
@@ C.2.2 Multi-agent Experiment
C.2.3 Human Validation of LLM Annotations
To validate the reliability of our LLM-assisted metrics, we conducted a human evaluation experiment to assess the accuracy of the annotations generated by Claude 3.7 Sonnet. Our research team manually annotated 5% of all negotiation messages across all six metrics. We then calculated the agreement rate between human annotations and LLM annotations to measure the reliability of our automated evaluation approach.
Table 11 presents the agreement rates for each negotiation metric. The results demonstrate high agreement between human and LLM annotations across all metrics, with accuracy rates ranging from 91.67% to 100%. This strong agreement suggests that the LLM-based annotation approach provides reliable evaluations for analyzing negotiation behaviors in Diplomacy.
| Negotiation Metric | Agreement Rate (%) |
| Reasoning and Negotiation Alignment | 98.51 |
| Mutual Benefit or Exploitative Proposal Nature | 97.01 |
| Peace or Conflict Messaging | 100.00 |
| Perspective Taking or Empathy | 91.04 |
| Conditional Plans | 97.01 |
| Proposal Acceptance Rate | 95.31 |
Appendix D Prompt Design and Interaction Flow
D.1 PDDL
D.2 Cooperative Game: Hanabi
@@
D.3 Strategic Game: Diplomacy
@@
Prompts for strategy alignment, mutual vs. one-sided benefits, peace vs. conflict messaging, perspective-taking and conditional plans:
Prompts for proposal acceptance:
Example Annotation by Claude 3.7 Sonnet: